Status Report

ISS On-Orbit Status 23 Aug 2002

By SpaceRef Editor
August 23, 2002
Filed under , ,

All ISS systems continue to function nominally, except as noted previously or below.

Crew activities today focused on preparations for the upcoming spacewalk.

Hatch open for EVA-8 is set for 1:02 am EDT Monday morning (still in eclipse), and hatch close for about 6:56am. During DC-1 airlock depress, station attitude will be put on Hold, and the SM’s #2 solar wing will be feathered just before egress.  Attitude thrusters will be inhibited during the spacewalk.  After the EVA, the ISS will be repressurized by O2 from Progress 8P and/or N2 (nitrogen) from the USOS Airlock to keep total atmospheric pressure above flight rule (FR) limit.  The EVA will be covered on NASA TV.

FE-2 Sergey Treschev began work last night by performing pressure checks on the SM BK-3 oxygen (O2) tanks and the BNP portable repress O2 tank in the DC-1 airlock module, as CDR Valery Korzun worked on his Orlan-M suit (#14) to deactivate a low-frequency C&W (caution and warning) tone.  [During EVA-7, this tone, generated by the primary subset of the Corona-M radio system in the Orlan’s BRTA telemetry system, interfered with his communications.  At the same time, Korzun tested the proper performance of the C&W tone generator in the Corona-M’s backup set.]

Korzun and Treschev then readied their suit ORUs (orbit replaceable units), such as LiOH (lithium hydroxide) canisters for CO2 (carbon dioxide) absorption, BK-3 O2 tanks, moisture collectors, feedwater filters, batteries, etc.  They also prepared personal gear (gloves, socks, undergarment, etc.), and Treschev adjusted the sizing straps of his Orlan spacesuit (#12) to his body dimensions.

Later in the day, both EV crewmembers installed the external EVA equipment on their suits and tagged up with a TsUP EVA specialist via S-band.

They also checked out their electrocardiogram harnesses (PKO-BETA08) on the GAMMA-1 biomedical examination panel (PKO) before they set up for a thorough test of the Orlan BRTA telemetry and communications from the DC-1.  Other preparations involved leak checks and valve function tests on the pressure suits and their support systems (BSS), as well as actuation verification of pressure equalization valves (KBD) from the EVA support panels (POV) in the PkhO and DC-1 modules.

FE-1 Peggy Whitson closed out the EMU (US extravehicular mobility unit) battery discharge operations from yesterday and initiated the recharge of the batteries, to go on for an estimated 28 hours.  [After termination tomorrow, the crew should have two good EMU batteries in case a contingency EMU EVA should become necessary.]

Whitson connected the UOP (utility operations panel) bypass cable to the Lab RWS DCP (robotic workstation/display & control panel), in preparing for Monday’s use of the SSRMS video cameras to cover the EVA.

Peggy Whitson was scheduled to set up and activate the GASMAP (gas analyzer system for metabolic analysis physiology) in the HRF (human research facility) and the PuFF (pulmonary function in flight) equipment and then to calibrate the equipment for Valery Korzun to take a 40-min. standard pre-EVA test.  [PuFF investigates the effects of EVA and long-term exposure to zero-G on the pulmonary (lung) function.  It utilizes the GASMAP along with other equipment, such as disposable mouthpiece, manual breathing valve, flowmeter, pressure-flow module, pressure and volume calibration syringes.  Each PuFF session, occurring monthly, involves five lung function tests.]

Regular daily routine maintenance was performed by Whitson (SOSh life support systems servicing and Lab payload status checkup) and Treschev (IMS delta file preparation).

All crewmembers performed their daily physical exercise program on TVIS treadmill, RED expander and VELO cycle with load trainer.

Troubleshooting activity is being developed for a TVIS comm problem that currently prevents the transfer of exercise data from the treadmill’s PCMCIA (personal computer memory card international association) to the MEC (medical equipment computer) for subsequent downlink to Earth.

Last Wednesday’s (8/21) inspection of the TVIS uncovered a second frozen truss roller, requiring the crew to repeat the earlier procedure to remove the roller axle.  Meanwhile, the necessary hardware for fixing the damaged TVIS chassis is ready for shipment to Moscow and thence to Baikonur for loading on the next Progress, 9P, scheduled for launch on 9/20.

Last night, the ISS was maneuvered from inertial XPOP attitude (x-axis perpendicular to orbit plane) back to the thermally more friendly earth-oriented LVLH (local vertical/local horizontal) attitude. Since EVA-8 is planned to be done in LVLH anyway, the attitude change was moved forward due to concerns about CMG (control moment gyroscope) bearing temperatures.  After last night’s maneuver, bearing temperatures were seen to drop to more acceptable levels.  The station will go back to XPOP after Monday’s EVA, on Daily Orbit 7. [Analysis of detrimental effects of the thermal gradient between the inside and outside of the CMG spin bearings is continuing, to determine an acceptable upper temperature limit for XPOP periods.  XPOP is necessary at some solar Beta angles for photovoltaic power reasons, and there are also contrasting interests of Moscow, such as Progress overheating and payload pointing.  Thus, US and Russian specialists are hard at work to develop a best-compromise schedule for switching between the two flight attitudes in the future.]

The attitude issue is emphasized by bearing problems encountered during ground preparations of the spare CMG for its intended flight on ULF-1/STS-114 (NET 3/1/03).  While undergoing standard spin testing, “chirping” sounds from the CMG were noticed, which yesterday turned into a “metal-on-metal whine”.  The CMG was spun down with what is probably a bearing failure.  It will be investigated, and a plan will be developed for getting the spare ready.   The issue heightens the concern about the on-orbit CMGs’ bearings and underlines the importance for including their consideration in XPOP vs. LVLH scheduling.

Inspection by the crew of the trash bag containing the leaky CWC (collapsible water container (#5087) showed that its water has leaked into the trash bag.  MCC-H has determined that the CWC cannot be used for condensate processing due to concerns for microbial contamination.  Evaluations are going on to determine how to soak up the water and package bag and all for disposing.

Troubleshooting of the MCA (major constituent analyzer) has been completed, and several good workaround options have been developed to bring it back.

Tests are showing that the pressure reduction process inside battery 4B2 of the P6 module’s 4B power channel is proceeding satisfactorily. The activity should be finished in one to two days.

Huntsville’s POC (Payload Operations Center) has completed development of a clean-up procedure for the MSG (Microgravity Science Glovebox) with the fractured SUBSA ampoule.  The task is scheduled for the crew on 8/30 (Friday).

278 kg of propellants have been transferred from Progress 8P to the SM, leaving 432 kg in the cargo ship.

Future launches:  New date for Mission 9A/STS-112 is now 10/2.  Mission 11A/STS-113 will launch on or immediately after the day of Soyuz 5S undocking (11/6).  

Today’s targets for the CEO (crew earth observations) program were Taiwan Smog (over dense W coast urban-industrial zone of the island, on out into the sea towards the mainland coast. Crew was to shoot oblique views for this smoggiest part of the planet), Beijing, China (nadir view.  Mostly sunny after weeks of cloudy weather. ESC [electronic still camera] requested), Tianjin, China (Beijing‚s vast port complex [slightly right of nadir, SE Beijing]. ESC requested), Tel Aviv, Israel (nadir view.  Major coastal city. ESC requested), Damascus, Syria (nadir view.  Of interest: to look on the dry side of the Lebanese mountains.  This city is more difficult to see than cities on the greener mountains and east coast.  ESC requested), and Eastern Mediterranean Aerosols (high pressure building in.  Crew was to photograph any aerosols concentrated in the narrow Adriatic Sea [between confining landmasses-looking left], and in the eastern Med [looking right].  Smog from N Italy gets to Athens and Turkey regularly. Second pass offers good opportunity to shoot Europe‚s most consistently smoggy region, the Po River valley of northern Italy).
CEO images can be viewed at the website

U.S. and Russian Segment Status (as of 10:42 am EDT):

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is powered On (32-amp mode), on backup pump.  Vozdukh CO2 scrubber is ON in MANUAL cycle mode #5, i.e., 10-min. cycle time (vacuum pump failed).  U.S. CDRA CO2 scrubber is on Override.  BMP Harmful Impurities unit: Absorbent bed #1 in Purify mode, bed #2 in Purify mode.
  • SM Working Compartment:  Pressure (mmHg) — 758, temperature (deg C) — 27.4, ppO2 (mmHg) — 157.5, ppCO2 (mmHg) — 2.6.
  • SM Transfer Compartment:  Pressure (mmHg) — 751, temperature (deg C) — 20.5.
  • FGB Cabin:  Pressure (mmHg) — 756, temperature (deg C) — 20.0.
  • Node:  Pressure (mmHg) — 752.16, temperature (deg C) — 23.9 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • U.S. Lab:  Pressure (mmHg) — 753.91, temperature (deg C) — 24.7, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a;
  • Joint Airlock (Equip. Lock):  Pressure (mmHg) — 753.91, temperature (deg C) — 27.6; shell heater temp (deg C) — 27.4, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • PMA-1:  Shell heater temp (deg C) — 22.9
  • PMA-2:  Shell heater temp (deg C) — 16.8

(Note: Partial pressures ppO2 and ppCO2 in U.S. segment [USOS] not available because MCA [major constituent analyzer] is failed and in Extended Life mode [= a state that preserves mass spectrometer vacuum but produces no pp data]). MSA (mass spectrometer assembly) and VGA (verification gas assembly) were replaced, but some more work needs to be done).

Electrical Power Systems (EPS):

  • Both P6 channels fully operational.  Beta Gimbal Assembly (BGA) 2B and BGA 4B in Autotrack mode (solar-tracking).
  • SM batteries:   Battery #1 is off-line; all other batteries (7) are in “Partial Charge” mode.
  • FGB batteries:  Battery #5 is off-line, battery #6 is in “Cycle” mode; all other batteries (4) are in “Partial Charge” mode.
  • Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby mode.

Thermal Control Systems:

  • Air conditioner SKV-1 is On. SKV-2 is Off.

Command & Data Handling Systems:

  • C&C-3 MDM is prime, C&C-2 is back-up, and C&C-1 is in standby.
  • GNC-1 MDM is prime; GNC-2 is Backup.
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-1 MDM is operational; PL-2 MDM is Off.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 3 redundant lanes (of 3) operational.
  • SM Central Computer (TsVM): 3 redundant lanes (of 3) operational.

Attitude Source:

  • 3 CMGs on-line.
  • State vector — US GPS (SIGI string 1)
  • Attitude — Russian segment
  • Angular rates — US RGA1 (rate gyro assembly 1)

Communications & Tracking Systems:

  • All Russian communications & tracking systems are nominal.
  • S-band is operating nominally.
  • Ku-band is operating nominally.
  • Audio subsystem operating nominally.
  • Video subsystem operating nominally.
  • MCOR (medium-rate communications outage recorder) is operating nominally


  • SSRMS/Canadarm2 at MBS PDGF 1 (mobile base system/power & data grapple fixture 1), in EVA-7 viewing position, with Keep Alive power on both strings (based on MBS).
  • MBS: Keep Alive power on both strings.
  • RWS (robotics workstations): Lab RWS is Off; Cupola RWS is Off.

ISS Orbit (as of this morning, 3:56am EDT [= epoch]):

  • Mean altitude — 392.4 km
  • Apogee — 403.8 km
  • Perigee — 381.0 km
  • Period — 92.4 min.
  • Inclination (to Equator) — 51.64 deg
  • Eccentricity — 0.0016816
  • Orbits per 24-hr. day — 15.58
  • Altitude decrease — 250 m (mean) in last 24 hours
  • Solar Beta Angle — -38.8 deg (magnitude increasing)
  • Revolutions since FGB/Zarya launch (Nov. ‚98) — 21449
  • Current Flight Attitude — LVLH (local vertical/local horizontal = „earth-fixed‰: z-axis in local vertical, x-axis in velocity vector [yaw: -10 deg, pitch: 7.8 deg, roll: 0 deg]).  Next change will be back to XPOP on 8/26.

For more on ISS orbit and worldwide naked-eye visibility dates/times, see

SpaceRef staff editor.